### v2

@@ Abstract
Recently, Large Language Models (LLMs) have dominated much of the artificial intelligence scene with their ability to process and generate natural languages. However, the majority of LLM research and development remains English-centric, leaving low-resource languages such as those in the Southeast Asian (SEA) region under-represented. To address this representation gap, we introduce Llama-SEA-LION-v3-8B-IT and Gemma-SEA-LION-v3-9B-IT, two cutting-edge multilingual LLMs designed for SEA languages. The SEA-LION family of LLMs supports 11 SEA languages, namely English, Chinese, Indonesian, Vietnamese, Malay, Thai, Burmese, Lao, Filipino, Tamil, and Khmer. Our work leverages large-scale multilingual continued pre-training with a comprehensive post-training regime involving multiple stages of instruction fine-tuning, alignment, and model merging. Evaluation results on multilingual benchmarks indicate that our models achieve state-of-the-art performance across LLMs supporting SEA languages. We open-source the models [^1] SEA-LION Models Collection to benefit the wider SEA community.
@@ 1 Introduction
Recently, there have been many attempts to create multilingual LLMs in an open-source manner, e.g., BLOOM (Scao et al., 2022), a project aimed at increasing multilingual presence in open-source LLMs by supporting 46 natural languages. Popular LLM families such as Llama (Dubey et al., 2024), Gemma (Rivière et al., 2024) and Qwen (Yang et al., 2024a) have also introduced multilingual LLMs for their latest iteration. During our evaluations, we found that the performance of these models is acceptable in the general case, i.e., if we consider evaluation benchmarks formulated from English datasets, but we observe that the performance degrades on SEA-specific benchmarks. Moreover, researchers have also introduced LLMs such as SeaLLMs Nguyen et al. (2024); Zhang et al. (2024a) and Sailor Dou et al. (2024) to specifically address the LLM gap in SEA languages. However, the performance of these models is less than ideal for languages such as Thai or Tamil [^2] Tamil is one of the official languages in Singapore. It is also spoken in other areas in the SEA region, such as Malaysia. (10X et al., 2024; AI Products Team, 2024).
In this paper, we address the issues by proposing a robust open-source Southeast Asian model with data transparency for reproducibility, namely SEA-LION – a family of LLMs CPT and fine-tuned on Llama-3.1-8B-Instruct and Gemma-2-9B with a focus on SEA languages. To tackle the performance problem, we utilize 200 billion English, code and SEA languages tokens as well as 16.8 million English and SEA languages instruction and answer pairs for CPT and post-training steps respectively, to achieve a significant improvement in SEA languages. In order to allow our models to be used by everyone without restrictions, we release our models under a fully open MIT license. We benchmark our models against the SEA-HELM [^3] SEA-HELM Leaderboard (Susanto et al., 2025) and Open LLM Leaderboard [^4] Open LLM Leaderboard with other LLMs of similar sizes in Southeast Asia like Sailor 2 (Team, 2024) and SeaLLMs 3 (Zhang et al., 2024a) where our models achieve state-of-the-art performances. We summarize the contribution of our paper as follows.
    We released two LLMs, Llama-SEA-LION-v3-8B-IT and Gemma-SEA-LION-v3-9B-IT, that are meticulously trained to accurately represent the unique linguistic diversity of SEA languages.
@@ 2.1 Pre-training data
The CPT data consists of a curated set of English, multilingual and code corpora from several open source repositories like Dolma (Soldaini et al., 2024), FineWeb (Penedo et al., 2024), the-stackv2 (Lozhkov et al., 2024), SEA-LION-Pile (AI Singapore, 2023), SEA-LION-Pile-v2 (AI Singapore, 2025), as well as documents from CommonCrawl (CommonCrawl, 2024) and from the public domain such as Wikipedia (Foundation, 2024). For SEA-LION-Pilev2, we filter CommonCrawl WARC data for documents in SEA languages (i.e., Burmese, Simplified Chinese, Indonesian, Khmer, Lao, Malay, Filipino, Tamil, Thai and Vietnamese) using the pretrained fasttext language classifier (Joulin et al., 2017).
A document is retained if the language code reported in its metadata matches with that of one of the aforementioned SEA languages. Additionally, we further clean up the data with Trafilatura (Barbaresi, 2021). To determine the best dataset ratio between SEA languages, code and English for the CPT process, we perform a series of small-scale CPT experiments each with a training budget of 10B tokens and varying proportions of English, code and SEA language data. We settled on an optimal data mix ratio of 55% SEA languages, 25% English and 20% code tokens for a budget of 200B tokens. For a detailed breakdown of the token count by languages, please refer to the model card. [^5] Gemma-SEA-LION-v3-9B Model Card
@@ 2.2 CPT process
Training setup. Following previous works Dou et al. (2024), we use BPE-Dropout (Provilkov et al., 2020) to increase the performance and robustness of the training. We use a Warmup-Stable-Decay (WSD) (Hu et al., 2024) scheduler with warm-up and cooldown phases each representing 10% of the entire training budget. We use the AdamW (Loshchilov and Hutter, 2019) optimiser with the maximum learning rate (LR) set to ${1e^{-5}}$ and the final LR after cooldown is ${1e^{-7}}$. Following Wortsman et al. (2024), we set epsilon to ${1e^{-15}}$. We use Composer (Team, 2021) and LLM Foundry (Team, 2022) for distributed training using Fully Sharded Data Parallel (Zhao et al., 2023) on a cluster of eight nodes of the p5.48xlarge instance from Amazon Web Services (AWS). The total training duration was approximately 6 days and 10 days for the Llama 3.1 and Gemma 2 models, respectively. In this paper, we refer to the post-CPT models as Llama-SEA-LION-v3-8B and Gemma-SEA-LION-v3-9B for the Llama 3.1 and Gemma 2 continued pre-trained models respectively.
@@ 3.1 Post-training data
The post-training data for instruction fine-tuning consists of Infinity-Instruct [Foundation and Chat] (Beijing Academy of Artificial Intelligence, 2024), OpenMath-Instruct 2 (Toshniwal et al., 2024) and our own SEA-Instruct.
In particular, SEA-Instruct consists of multiple open-source instruction datasets, a synthetically generated dataset following the Magpie (Xu et al., 2024) template, and hand-crafted datasets collected from native Southeast Asians. The full detail of the SEA-Instruct and SEA-Preference dataset can be found in the model card. [^6] Gemma-SEA-LION-v3-9B-IT Model Card
Figure 1: Training process of Llama-SEA-LION-v3-8B-IT (Section 3.2.1). The post-training process consists of 2 stages of instruction fine-tuning, an alignment stage and multiple merge stages. Dotted lines denote a merge stage and solid lines denote an alignment stage.
@@ 3.2 Post-training process
3.2.1 Llama-SEA-LION-v3-8B-IT
Stage 1 IFT As shown in Figure 1, we started off the post-training phase with IFT of Llama-SEA-LION-v3-8B with the Infinity Instruct (Foundation) (Beijing Academy of Artificial Intelligence, 2024) and OpenMathInstruct2 Toshniwal et al. (2024) datasets. Both datasets contain approximately 9.5 million instruction pairs, primarily in English and centered around reasoning, math, and code. We refer to the model at this stage as Stage-1-Llama.
@@ 3.2.1 Llama-SEA-LION-v3-8B-IT
First merge After finishing the IFT stages, we performed the first of a series of merges by merging Stage-1-Llama and Stage-2-Llama into the Llama-SEA-LION-v3-8B using the DARE TIES (Yu et al., 2024; Ilharco et al., 2023) method. We refer to the model at this stage as Merge-1-Llama.
@@
Final merge Lastly, we perform a merge using the DELLA-Linear merge. With the original Llama-3.1-8B-Instruct model as the base for merging, we merge in Merge-2-Llama and Aligned-SimPO-Llama to produce the final model, Llama-SEA-LION-v3-9B-IT.
3.2.2 Gemma-SEA-LION-v3-9B-IT
Figure 2: Training process of Gemma-SEA-LION-v3-9B-IT (Section 3.2.2). The post-training process comprises two stages of instruction fine-tuning, an alignment stage, and multiple merge stages. Dotted lines denote a merge stage and solid lines denote an alignment stage.
Stage 1 and Stage 2 IFT Similar to the Llama-SEA-LION-v3-8B-IT, we started off the post-training phase with both stages of IFT using the same datasets on the Gemma-2-9B model (Rivière et al., 2024). We refer to both models at stage 1 and stage 2 as Stage-1-Gemma and Stage-2-Gemma, respectively.
@@ 3.2.2 Gemma-SEA-LION-v3-9B-IT
Final merge Finally, using the Gemma-2-9B model as the base model, we merged Merge-1-Gemma, FuseChat Gemma-2-9B-Instruct (Yang et al., 2024b), Gemma-SEA-LION-v3-9B, and Aligned-SimPO-Gemma into it to produce the final model Gemma-SEA-LION-v3-9B-IT.
@@ 3.3 Discussion
The entire post-training process for Gemma-SEA-LION-v3-9B-IT and Llama-SEA-LION-v3-8B-IT took approximately 1350 and 1024 GPU hours respectively on eight H100 GPUs. To make the training efficient, all post-training steps utilize Liger Kernel (Hsu et al., 2024) for substantial memory savings of approximately 60%.
4 Experimental Setup and Results
@@
4.1 Setup
For the evaluation, we compared our models against well-known LLMs: SeaLLMsv3 Zhang et al. (2024a), Sailorv2 Team (2024), Qwen 2.5 Yang et al. (2024a), Gemma 2 Rivière et al. (2024) and Llama 3.1 Dubey et al. (2024), where the parameters of those models are less than 10 billion parameters, similar to our models. The evaluations are split into two areas as follows.
Multilingual performance. We evaluated the multilingual performance of each LLM using the SEA-HELM Leaderboard (Leong et al., 2023; Susanto et al., 2025). Due to the lack of proper benchmarks for low-resource languages (e.g. Lao, Khmer, Filipino), we only benchmarked the languages covered in the SEA-HELM leaderboard, which are Indonesian, Tamil, Thai, and Vietnames [^7] SEA-HELM has been updated since this paper was written to include Filipino. We selected SEA-HELM because the design choice of this benchmark reflects the performance of SEA culture and knowledge the most compared with other existing benchmarks DAMO-NLP-SG (2024); Lovenia et al. (2024); Wang et al. (2024a). We used the evaluation code from the official website [^8] SEA-HELM Repository without any changes.
English performance. We evaluated the English performance of the models using the Open LLM Leaderboard (HuggingFace, 2024). The leaderboard consists of six benchmarks, IFEval Zhou et al. (2023), Big Bench Hard (Suzgun et al., 2023), MATH (Hendrycks et al., 2021), GPQA (Rein et al., 2023), MuSR (Sprague et al., 2024) and MMLU-PRO (Wang et al., 2024c).
4.2 Results

### v3

@@ Abstract
Recently, Large Language Models (LLMs) have dominated much of the artificial intelligence scene with their ability to process and generate natural languages. However, the majority of LLM research and development remains English-centric, leaving low-resource languages such as those in the SouthEast Asian (SEA) region under-represented. To address this representation gap, we introduce Llama-SEA-LION-8B-IT and Gemma-SEA-LION-9B-IT, two cutting-edge multilingual LLMs designed for SEA languages. The SEA-LION family of LLMs supports 11 SEA languages, namely English, Chinese, Indonesian, Vietnamese, Malay, Thai, Burmese, Lao, Filipino, Tamil, and Khmer. Our work leverages large-scale multilingual continued pre-training with a comprehensive post-training regime involving multiple stages of instruction fine-tuning, alignment, and model merging. Evaluation results on multilingual benchmarks show that our models achieve state-of-the-art performance across LLMs supporting SEA languages. We open-source the models111[LINK_HIDDEN] to benefit the wider SEA community.
@@ 1 Introduction
Recently, there have been many attempts to create multilingual LLMs in an open-source manner, e.g., BLOOM (Scao et al., 2022), a project aimed at increasing multilingual presence in open-source LLMs by supporting 46 languages. Popular LLM families such as Llama (Dubey et al., 2024), Gemma (Rivière et al., 2024) and Qwen (Yang et al., 2024a) have also introduced multilingual LLMs for their latest iteration. During our evaluations, we found that the performance of these models is acceptable in the general case, i.e., if we consider evaluation benchmarks formulated from English datasets, but we observe that the performance degrades on SEA-specific benchmarks. Moreover, researchers have also introduced LLMs such as SeaLLMs Nguyen et al. (2024); Zhang et al. (2024a) and Sailor Dou et al. (2024) to specifically address the LLM gap in SEA languages. However, the performance of these models is less than ideal for languages such as Thai or Tamil [^2] Tamil is one of the official languages in Singapore. It is also spoken in other areas in the SEA region, such as Malaysia. (10X et al., 2024; AI Products Team, 2024).
In this paper, we address the issues by proposing a robust open-source Southeast Asian model with data transparency for reproducibility, namely SEA-LION – a family of LLMs continued pre-trained (CPT) and fine-tuned on Llama-3.1-8B-Instruct for Llama-SEA-LION-8B-IT and Gemma-2-9B for Gemma-SEA-LION-9B-IT with a focus on SEA languages. To tackle the performance problem, we utilize 200 billion English, code and SEA languages tokens as well as 16.8 million English and SEA languages instruction and answer pairs for CPT and post-training steps respectively, to achieve a significant improvement in SEA languages. In order to allow our models to be used by everyone without restrictions, we release our models under a fully open MIT  [^3] https://mit-license.org/. We benchmark our models against the SEA-HELM(Susanto et al., 2025) and Open LLM Leaderboard [^4] Open LLM Leaderboard with other LLMs of similar sizes in Southeast Asia like Sailor 2 (Team, 2024) and SeaLLMs 3 (Zhang et al., 2024a), where our models achieve state-of-the-art performances. We summarize the contribution of our paper as follows.
    We released two LLMs, Llama-SEA-LION-8B-IT and Gemma-SEA-LION-9B-IT, that are meticulously trained to accurately represent the unique linguistic diversity of SEA languages.
@@
    We present a reproducible dataset development process, covering sourcing and the model training process. Upon the acceptance of this paper, we will release our training artifacts, including the training dataset, training scripts, training checkpoints, and fine-tuned models, including Llama-SEA-LION-8B-IT and Gemma-SEA-LION-9B-IT, to provide strong baselines, promote reproducibility, and enable future research on applications that require SEA-specific knowledge.
@@ 2.1 Pre-training data
The CPT data consists of a curated set of English, multilingual, and code corpora from several open source repositories like Dolma (Soldaini et al., 2024), FineWeb (Penedo et al., 2024), the-stackv2 (Lozhkov et al., 2024), SEA-LION-Pile (AI Singapore, 2023), SEA-LION-Pile-v2 (AI Singapore, 2025), as well as documents from CommonCrawl (CommonCrawl, 2024) and from the public domain, such as Wikipedia (Foundation, 2024). For SEA-LION-Pilev2, we filter CommonCrawl WARC data for documents in SEA languages (i.e., Burmese, Simplified Chinese, Indonesian, Khmer, Lao, Malay, Filipino, Tamil, Thai, and Vietnamese) using the pretrained fasttext language classifier (Joulin et al., 2017).
A document is retained if the language code reported in its metadata matches with that of one of the aforementioned SEA languages. Additionally, we further clean up the data with Trafilatura (Barbaresi, 2021). To determine the best dataset ratio between SEA languages, code and English for the CPT process, we perform a series of small-scale CPT experiments each with a training budget of 10B tokens and varying proportions of English, code and SEA language data. We settled on an optimal data mix ratio of 55% SEA languages, 25% English, and 20% code tokens for a budget of 200B tokens. For a detailed breakdown of the token count by languages, please refer to Table 6.
@@ 2.2 CPT process
Training setup. Following previous works Dou et al. (2024), we use BPE-Dropout (Provilkov et al., 2020) to increase the performance and robustness of the training. We use a Warmup-Stable-Decay (WSD) (Hu et al., 2024) scheduler with warm-up and cooldown phases each representing 10% of the entire training budget. We use the AdamW (Loshchilov and Hutter, 2019) optimiser with the maximum learning rate (LR) set to $1e^{-5}$} and the final LR after cooldown is $1e^{-7}$}. Following Wortsman et al. (2024), we set epsilon to $1e^{-15}$}. We use Composer (Team, 2021) and LLM Foundry (Team, 2022) for distributed training using Fully Sharded Data Parallel (Zhao et al., 2023) on a cluster of eight nodes of the p5.48xlarge instance from Amazon Web Services (AWS). The total training duration was approximately 6 days and 10 days for the Llama 3.1 and Gemma 2 models, respectively. In this paper, we refer to the post-CPT models as Llama-SEA-LION-8B and Gemma-SEA-LION-9B for the Llama 3.1 and Gemma 2 continued pre-trained models respectively.
@@ 3.1 Post-training data
The post-training data consists of 3 subsets of data for Stage 1 IFT, Stage 2 IFT, and the Preference dataset for alignment, respectively. We describe the training data information of each step as follows.
Stage 1 IFT. In this step, we employ Infinity-Instruct [Foundation and Chat] (Beijing Academy of Artificial Intelligence, 2024) and OpenMath-Instruct 2 (Toshniwal et al., 2024) to improve the mathematical, reasoning, and coding skills of the instruction model. The full details of the training data are shown in Appendix 7.
Stage 2 IFT. Then, in this step, we use generalized large-scale instructions on the previous instruction model. In particular, we employ 22 existing datasets (written in English, Thai, and Vietnamese) and formulate new 22 synthetic datasets using various models and techniques to create SEA instruction datasets (see Appendix A.3 for the full data generation details). As shown in Appendix 9, we use a total of 7,298,828 instruction samples that cover 11 languages.
Helpfulness and preference alignment. We also conduct an alignment learning on top of the instruction model using a feedback dataset called UltraFeedBack Cui et al. (2024). In addition, we also synthesized the SEA version of the UltraFeedBack using NemoTron-70b with Gemma2 as a reward model, see Appendix A.4 for the full details.
Figure 1: Training process of Llama-SEA-LION-8B-IT (Section 3.2.1). The post-training process consists of 2 stages of instruction fine-tuning, an alignment stage and multiple merge stages. Dotted lines denote a merge stage and solid lines denote an alignment stage.
@@ 3.2 Post-training process
3.2.1 Llama-SEA-LION-8B-IT
Stage 1 IFT As shown in Figure 1, we started off the post-training phase with IFT of Llama-SEA-LION-8B with the Infinity Instruct (Foundation) (Beijing Academy of Artificial Intelligence, 2024) and OpenMathInstruct2 Toshniwal et al. (2024) datasets. Both datasets contain approximately 9.5 million instruction pairs, primarily in English and centered around reasoning, math, and code. We refer to the model at this stage as Stage-1-Llama.
@@ 3.2.1 Llama-SEA-LION-8B-IT
First merge After finishing the IFT stages, we performed the first of a series of merges by merging Stage-1-Llama and Stage-2-Llama into the Llama-SEA-LION-8B using the DARE TIES (Yu et al., 2024; Ilharco et al., 2023) method. We refer to the model at this stage as Merge-1-Llama.
@@
Final merge Lastly, we perform a merge using the DELLA-Linear merge. With the original Llama-3.1-8B-Instruct model as the base for merging, we merge in Merge-2-Llama and Aligned-SimPO-Llama to produce the final model, Llama-SEA-LION-9B-IT.
3.2.2 Gemma-SEA-LION-9B-IT
Figure 2: Training process of Gemma-SEA-LION-9B-IT (Section 3.2.2). The post-training process comprises two stages of instruction fine-tuning, an alignment stage, and multiple merge stages. Dotted lines denote a merge stage and solid lines denote an alignment stage.
Stage 1 and Stage 2 IFT Similar to the Llama-SEA-LION-8B-IT, we started off the post-training phase with both stages of IFT using the same datasets on the Gemma-2-9B model (Rivière et al., 2024). We refer to both models at stage 1 and stage 2 as Stage-1-Gemma and Stage-2-Gemma, respectively.
@@ 3.2.2 Gemma-SEA-LION-9B-IT
Final merge Finally, using the Gemma-2-9B model as the base model, we merged Merge-1-Gemma, FuseChat Gemma-2-9B-Instruct (Yang et al., 2024b), Gemma-SEA-LION-9B, and Aligned-SimPO-Gemma into it to produce the final model Gemma-SEA-LION-9B-IT.
@@ 3.3 Discussion
The entire post-training process for Gemma-SEA-LION-9B-IT and Llama-SEA-LION-8B-IT took approximately 1350 and 1024 GPU hours, respectively, on eight H100 GPUs. To make the training efficient, all post-training steps utilize Liger Kernel (Hsu et al., 2024) for substantial memory savings of approximately 60%.
4 Experimental Setup
4.1 Competitive methods
For the evaluation, we compared our models against well-known LLMs for multilingual and SEA languages, such as SeaLLMsv3 Zhang et al. (2024a), Sailorv2 Team (2024), Qwen 2.5 Yang et al. (2024a), Gemma 2 Rivière et al. (2024) and Llama 3.1 Dubey et al. (2024), where the parameters of those models are less than 10 billion parameters, similar to our models.
SEA-HELM
NLU, NLG, NLR, NLI Instruction Following
Models Average ID VI TH TA ID VI TH
Meta-Llama-3.1-8B 35.37 42.33 40.67 35.13 38.88 16.19 19.05 9.00
SeaLLMs-v3-7B 37.04 44.79 48.29 43.53 27.45 26.67 35.24 26.00
Gemma-2-9B 41.48 47.65 43.28 42.00 53.26 4.76 3.81 10.00
Qwen2.5-7B 41.98 51.63 52.17 46.55 36.60 31.43 36.19 30.00
Sailor2-8B 42.62 53.23 47.33 46.64 45.04 30.48 30.48 35.00
Llama-SEA-LION-8B 41.42 44.98 46.25 42.79 43.03 25.71 32.38 23.00
Gemma-SEA-LION-9B 48.67 57.16 49.39 47.16 60.56 25.71 20.00 27.00
Open LLM Leaderboard
Models Average MMLU-PRO BBH GPQA MATH Lvl 5 IFEval (EN) MUSR
Meta-Llama-3.1-8B 13.9 24.95 25.29 6.32 5.14 12.7 8.98
Sailor2-8B 17.71 25.74 27.62 4.87 7.02 21.95 19.03
Gemma-2-9B 21.15 34.48 34.1 10.51 13.14 20.4 14.3
SeaLLMs-v3-7B 24.00 35.71 34.57 9.28 18.81 32.94 12.68
Qwen2.5-7B 24.99 37.39 35.81 9.96 18.88 33.74 14.14
Llama-SEA-LION-8B 16.61 27.6 26.04 7.49 9.89 16.56 12.07
Gemma-SEA-LION-9B 22.41 32.78 37.24 10.29 9.89 30.12 14.11
4.2 Evaluation Benchmarks
To evaluate the robustness of our proposed models, we compare our models to competitors in three benchmarks.
SEA Benchmarks. We evaluated the multilingual performance of each LLM using the SEA-HELM Leaderboard (Leong et al., 2023; Susanto et al., 2025). We selected SEA-HELM because the design choice of this benchmark reflects the performance of SEA culture and knowledge the most compared with other existing benchmarks DAMO-NLP-SG (2024); Lovenia et al. (2024); Wang et al. (2024a). We also evaluate on a wide-range SEA coverage language benchmark called SEACrowd Lovenia et al. (2024). This benchmark consists of all SEA languages for natural language understanding and generation datasets. However, due to maintenance reasons, we cannot reproduce the NLG benchmark of SEACrowd. Therefore, we experiment only with the NLU benchmark (zero-shot), which has 131 data subsets, 7 tasks, and 31 SEA indigenous languages.
English performance. We also evaluated the English performance of the models using the Open LLM Leaderboard (HuggingFace, 2024). This is because English is also widely used in SEA countries. Therefore, we need to evaluate the understanding and knowledge of LLMs in the English benchmark as well. The leaderboard consists of six benchmarks, IFEval Zhou et al. (2023), Big Bench Hard (Suzgun et al., 2023), MATH (Hendrycks et al., 2021), GPQA (Rein et al., 2023), MuSR (Sprague et al., 2024) and MMLU-PRO (Wang et al., 2024c). Moreover, we also evaluate the CPT models on SEA-HELM and the Open LLM Leaderboard since these benchmarks support the CPT evaluation.
5 Experimental Results
@@
Figure 3: Zero-shot model performance across NLU tasks in SEA languages.
Model NLU Score
SeaLLMs-v3-7B-chat 52.68
Llama-3.1-8B-Instruct 49.94
Sailor2-8B-Chat 60.21
Qwen2.5-7B-Instruct 54.51
Gemma-2-9B-IT 60.21
Llama-SEA-LION-8B-IT 55.10
Gemma-SEA-LION-9B-IT 64.13
Table 4: The average NLU performance across 131 data subsets and 31 indigenous languages.
@@
To understand the robustness and generalization of our proposed models, we conduct three studies as follows. Section 5.1 evaluates the robustness of continual pre-training models using SEA-HELM and the Open LLM leaderboard. In Section 5.2, we compare our instruction fine-tuning models with competitors in three benchmarks to demonstrate the generalization of our models. Lastly, we discuss the design choice of our models in Section 5.3.
5.1 Continued Pre-Training Results
SEA performance. The CPT stage is primarily focused on gaining SEA language capabilities and knowledge. For the purpose of comparison against base and CPT models, as shown in Table 1, we observed a 6.05 and 7.19 average SEA-HELM performance increase over the Meta-Llama-3.1-8B and Gemma-2-9B for Llama-SEA-LION-8B and Gemma-SEA-LION-9B, respectively. We observed a much larger average increase with instruction following capabilities in particular, which we attribute to the fact that our CPT models are trained from the instruction models rather than from the base models. Moreover, in the average performance, we found that our Gemma-SEA-LION-9B models perform the best compared to other models. This emphasizes a strong reason to perform CPT for improving the performance of SEA languages, rather than skipping the CPT and performing SFT directly.
English performance. For the English performance, as shown in Table 2, both CPT models also managed to perform competitively against the Meta-Llama-3.1-8B and Gemma-2-9B base models on the Open LLM Leaderboard benchmarks. This indicates that our choice of retraining with a proportion of 25% English tokens has been beneficial in mitigating catastrophic forgetting, which has been shown to stem from CPT (Zheng et al., 2024a). Although our CPT models perform lower than Qwen and SeaLLMs on this benchmark, we outperform them on the SEA language instead, which is the main focus of this work.
5.2 Instruction Fine-tuning Results
In this study, we compare our models with competitors on SEA-HELM, SEACrowd, and the Open LLM Leaderboard as follows.